The number of cars sold per day at a local car dealership, along with its corresponding probabilities, is shown in the table. Compute for the mean, variance, and standard deviation of the probability distribution of the random variable X. 


1
Expert's answer
2022-01-31T12:03:02-0500

Solution:

Assume the given table is:


"\begin{array}{c:c}
 x &Pr(X=x) \\
 \hdashline
0 & 0.1
\\ 
 \hdashline
1 & 0.1\\
 \hdashline
2 & 0.2\\
 \hdashline
3 & 0.3\\
 \hdashline
 4 & 0.2
\\
 \hdashline
 5 & 0.1
\end{array}"


Now, mean"=E(X)=\Sigma x.P(x)=0\times0.1+1\times0.1+2\times0.2+3\times0.3+4\times0.2+5\times0.1"

"=0.1+0.4+0.9+0.8+0.5
\\=2.7"

"E(X^2)=\Sigma x^2.P(x)"

"=0^2\times0.1+1^2\times0.1+2^2\times0.2+3^2\times0.3+4^2\times0.2+5^2\times0.1
\\=0.1+0.8+2.7+3.2+2.5
\\=9.3"

Then, "variance=E(X^2)-[E(x)]^2=9.3-2.7^2=2.01"

And standard deviation"=\sqrt{variance}=\sqrt{2.01}=1.417"